OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2)

OpenStudy (anonymous):

Do you know how to find the slope?

OpenStudy (anonymous):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]

OpenStudy (anonymous):

so m= 2-1/-4+6

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

Follow the steps :)

OpenStudy (anonymous):

Plug in (x1,y1) for point slope form, leave the x and y alone.

OpenStudy (anonymous):

so y-1=m(x+6) ??

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

or would i solve for M and then put it in y=mx+b form?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Looks good, but now make sure you plug in slope for m, because m = slope :)

OpenStudy (anonymous):

how

OpenStudy (anonymous):

y-1=m(x+6) you notice the m here, it's your slope the first thing we solved for

OpenStudy (anonymous):

man im tellin u i hate who ever makes us do slope in school lol

OpenStudy (anonymous):

Haha it's alright :), it's very useful in higher mathematics!

OpenStudy (anonymous):

okay so m= 1/2

OpenStudy (anonymous):

Looks good, now just make your equation look like y = mx+b form and we're done

OpenStudy (anonymous):

so 1=(1/2)-6+b

OpenStudy (anonymous):

\[y-1=1/2(x-6) \implies y - 1 = 1/2x-3\]

OpenStudy (anonymous):

how the crap did u get that lol

OpenStudy (anonymous):

\[y = 1/2x-3+1 \implies y = 1/2x+2\] and now we have equation of the line

OpenStudy (anonymous):

Follow the steps I gave you, you were doing good to :)

OpenStudy (anonymous):

so my equation is y=1/2x+2

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

It should be -2

OpenStudy (anonymous):

y=1/2x-2

OpenStudy (anonymous):

why didn't we just change x and y into numbers since we knew what they were. and how did u get b lol

OpenStudy (anonymous):

We didn't know what they were

OpenStudy (anonymous):

x1 is -6 y1 is 1

OpenStudy (anonymous):

\[(x_1, y_1) \neq (x,y)\]

OpenStudy (anonymous):

ugghh idk i hate math so much

OpenStudy (anonymous):

Just go over everything we did, all the information is here, I think you will figure it out

OpenStudy (anonymous):

i have been lookin over it the whole time haha still dont understand a thing

OpenStudy (anonymous):

Alright lets go over it together then

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

We have the points \[(-6,1) = (x_1,y_1) ~~~~ (-4,2) = (x_2,y_2)\] So these points will help us first find the slope, \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} } = \frac{ 2-1 }{ -4+6 } = \frac{ 1 }{ 2 }\] geometrically all a slope is saying is there is a change in coordinates, so something like this lets say